OpenStudy (omnomnom):

A object is thrown straight upwards with a speed of 19.6 m/sec. Three seconds later, its speed is: Select one: a. zero. b. 9.8 m/sec downward. c. 49 m/sec upward. Incorrect d. 29.4 m/sec downward.

OpenStudy (omnomnom):

@bahrom7893

OpenStudy (bahrom7893):

Vf = Vi+at = Vi+gt Vf = 19.6 + (-9.81)*3 = -9.83 => 9.8 down

OpenStudy (anonymous):

\[v=u+at\]

OpenStudy (anonymous):

Since it is thrown a=-g

OpenStudy (omnomnom):

wait i know its not cso would it be a?

OpenStudy (bahrom7893):

it's b

OpenStudy (omnomnom):

why?